# A stochastic optimization algorithm for revenue maximization in a service system with balking customers

### v1

@@ Abstract
SAB is with Korteweg-de Vries Institute for Mathematics, University of Amsterdam, Amsterdam, The Netherlands. HH is with School of Industrial Engineering, Purdue University, West-Lafayette IN, United States. MM is with the Mathematical Institute, Leiden University, The Netherlands; MM is also affiliated with (a) Korteweg-de Vries Institute for Mathematics, University of Amsterdam, Amsterdam, The Netherlands, b) Eurandom, Eindhoven University of Technology, Eindhoven, The Netherlands, (c) Amsterdam Business School, Faculty of Economics and Business, University of Amsterdam, Amsterdam, The Netherlands. LR is with Department of Statistics, University of Haifa, Israel.
@@ 1 Introduction
In this paper, we examine a service facility represented by a single-server queue, where customers arrive in response to the price set by the provider. The service provider’s objective is to determine the optimal price, i.e., the price that maximizes the expected revenue per unit of time. Achieving this goal is particularly challenging, as it involves a nuanced understanding of the intricate relationship between pricing, customer demand, system congestion, and the resulting impact on revenue performance.
Offline and offline revenue maximization — To put our work into perspective, we first provide a high-level categorization of the existing literature on revenue optimization. Broadly speaking, two dominant approaches can be identified, that one could refer to as the offline and online paradigms.
@@
    Offline revenue optimization is typically studied using a demand curve $d(p)$ that depends on the price $p$. The standard static optimization problem takes the form $\max_{p\geq 0}p\,d(p)$. However, in the context of service systems, an additional layer of complexity arises: congestion can also reduce demand. Specifically, users may decide to balk if the (expected) delay exceeds their individual utility thresholds. In stylized Markovian models, one can explicitly derive the revenue-maximizing price. A canonical example is Naor’s model [21], which features an observable queue and linear waiting costs. In such settings, the effective demand, captured by the effective arrival rate $\lambda(p)$, depends on both the price and the stationary queue-length distribution. The problem of monopolistic pricing in congested systems has been studied extensively under various modeling assumptions; we refer the reader to [7, 12] for comprehensive reviews. In studies where congestion is modeled through a queueing mechanism, the objective function often includes a deduction for queueing-related costs — see, for example, [17, 18], in which a term proportional to the mean stationary queue length is subtracted. These analyses typically assume that the primitives of the underlying queueing model (in particular the service-time distributions) are known.
@@
    In the dynamic pricing literature, the optimal price is iteratively adjusted until, with some level of confidence, it can be said that the revenue-maximizing price has been found. This line of work typically assumes uncertainty in the demand curve, which must be learned, and in addition that the queueing primitives are not, or partly, known. Such a setting naturally calls for online algorithms that learn and optimize simultaneously; we refer to, e.g., [23] for a survey of this strand of the literature. An added layer of complexity arises when customers are delay-sensitive — for instance, when they decide not to enter the system upon observing excessive congestion. In the queueing literature this is typically handled by considering stylized models. For example, in [1] it is shown that a Bayesian algorithm converges to the optimal price for a small-buffer Markovian system with two customer types.
@@
The settings considered come in various flavors. First, they may differ in the components included in the revenue function. While the core component reflects income from service, many formulations also include a penalty term for performance degradation — commonly modeled as a function of mean queueing delay. Second, there are differences in the definition of the decision space. In some studies, the objective is solely to determine the optimal price, while others also optimize over additional parameters such as the service rate or the number of servers. Third, and specific to the model-free approach, there are several variants based on the type of information available to the service provider when making pricing decisions (e.g., the customers’ arrival times, queueing delays, etc.).
@@
Both offline and online approaches have distinct strengths and limitations. We begin by outlining key considerations related to the offline setting. (1) In this setting, it is typically assumed that the revenue can be expressed as a function of price. While evaluating this objective is tractable in simple queueing models, often relying on exponential inter-arrival and service times, it becomes considerably more challenging when incorporating more realistic features, such as general service-time distributions or realistic customer behaviors like balking upon encountering excessive workloads at arrival. (2) The offline approach also requires knowledge of the queueing model’s primitives and the demand curve. While certain customer parameters can often be estimated from historical data (e.g., the service-time distribution), the demand curve is more difficult to infer, as it demands observations across a broad range of pricing scenarios. An important but often overlooked issue is model uncertainty: the effects of estimation errors are typically ignored. (3) When revenue cannot be expressed explicitly, the offline approach frequently resorts to heavy-traffic approximations, which involve scaling up arrival rates and service capacity. This can make the optimal price sensitive to system size, entailing that in contexts where such scaling is inappropriate, the accuracy of these approximations degrades.
@@
We now turn to the advantages and challenges of online pricing approaches. A key strength of these methods is that they do not require an explicit revenue function or prior parameter estimation. Instead, prices are adjusted dynamically based on observed system behavior. The central challenge lies in devising a price update mechanism. An natural strategy is to define cycles within which the price remains fixed, and based on the observed behavior adapt the price. This setting can be framed as a multi-armed bandit problem, where each arm corresponds to a different price, and the objective is to identify the price that maximizes expected revenue. However, unlike classical bandits with independent rewards, queueing systems exhibit memory effects: customers carried over from previous cycles influence the current system state and, consequently, the observed revenue. Therefore, any effective algorithm must account for these transient dynamics when evaluating performance at different price levels. Another challenge concerns how prices are updated. A common gradient-based method adjusts the price iteratively using an estimate of the objective’s gradient. The main difficulty is estimating this gradient accurately from data while ensuring fast convergence. The ultimate goal is to develop algorithms with low regret, meaning minimal cumulative revenue loss relative to the optimal fixed price over time.
Our setup — As mentioned, the class of revenue maximization problems we consider requires a sound modeling of the interaction between price, demand, and congestion. While other approaches typically separate these effects — combining a demand-based revenue term with a congestion penalty (often proportional to the mean queueing delay) — our framework integrates congestion effects directly into the demand-related revenue term.
@@
Our base model is a single-server system with Poisson arrivals and a general service-time distribution. Upon arrival, customers are informed of the current price $p$ set by the administrator and the current workload level $v$, and decide whether to join or balk. Let $H_{\theta}(p,v)$ denote the probability that a customer joins when observing workload $v$ and price $p$, where $\theta$ is a known (possible multidimensional) parameter characterizing the customer population. This setup is flexible in the sense that it can represent a heterogeneous population of strategic customers with some parametric utility distribution, or it can simply reflect a behavioral (e.g., empirical) joining rule. The objective of the administrator is to maximize long-term revenue, i.e., they aim to find $p^{*}$ that maximizes
@@
where $V\equiv V(p)$ is the stationary workload distribution at price $p$. Thus, the effective arrival rate in this dynamic pricing problem is
@@
This type of optimization problem naturally lends itself to a stochastic gradient descent (SGD) approach. However, implementing such an approach presents significant challenges: (1) the gradient of $\lambda(p)$ must first be estimated, and (2) the bias and variance structure of the resulting estimator must be understood to establish meaningful asymptotic guarantees for the algorithm’s performance. Throughout this work, we assume that $\theta$ is known, as it can be efficiently estimated in general settings with balking customers, as discussed in [4, 13].
@@
Online learning for control of queueing systems — Stochastic approximation (SA) algorithms, the class of iterative methods for finding the roots or extrema of functions when only noisy observations are available, constitute a flexible and robust approach for solving stochastic optimization problems [6]); a specific subclass is SGD, which relies on noisy gradient estimates. In the context of queueing systems there is a difficulty that stationary expectations and their gradients are often intractable, which makes direct implementation of standard SA algorithms unfeasible. A useful method to overcome this difficulty is infinitesimal perturbation analysis (IPA); see [11] for an extensive introduction to this method and [16] for an example of its use in a queueing context. The regenerative properties of such systems can often be used in order to obtain unbiased estimators of the objective function and to verify convergence to the true solution [10]. In [22] an SA algorithm based on busy periods is used to evaluate the Nash equilibrium for queues with balking customers, and more general strategic customer behavior. Reinforcement learning is another related approach for steady-state optimization of queueing networks [9], particularly well-suited to settings with a complex or high-dimensional policy space.
@@
Our work follows in the tradition of [8], in that it also develops an SGA-based online learning-based dynamic pricing method aimed at identifying the revenue-maximizing price in a setting that includes congestion effects. In [8], these congestion effects are incorporated into the objective function through a term proportional to the mean stationary queue length, while the arrival rate depends solely on the current price level. It is assumed that the function $\lambda(p)$ is known, while the customers’ service-time characteristics are not. The price adaptation scheme involves observations of customer delays, where in the proofs it plays a central role that the queue process is geometrically ergodic [8, Lemma 2]. This approach necessitates assuming that the moment-generating and cumulant-generating functions of service and inter-arrival times satisfy certain Cramér-type conditions. The analysis relies on a specific construction for the inter-arrival and service time distributions to enable the use of conventional IPA. Specifically, their setup involves a sequence of independent and identically distributed random variables $(U_{n})_{n\in\mathbb{N}}$, where the $n$-th inter-arrival time, for a given price $\lambda(p)$, is modeled as $U_{n}/\lambda(p)$; a similar construction is applied to the service times. Notably, this definition implies that the effective arrival process is not congestion-dependent. In contrast, our approach avoids imposing such a structure, in that we assume that the inter-arrival distribution depends on both the price and the workload.
@@
    Regarding the problem formulation, we explicitly model the behavior of clients who choose not to join the system when faced with high congestion, resulting in lost revenue. Incorporating such balking customers is particularly challenging, as in our framework we observe only the effective arrivals — that is, data related to the customers who do not balk and enter the system. The dependence of each customer’s joining decision on the congestion level introduces several evident technical hurdles. A useful insight is that the arrival rate at a given price $p$ can be written as the reciprocal of the mean effective inter-arrival time at the same price.
@@
    We develop an algorithm that provably converges to the optimal price $p^{*}$. The adaptive queue control schemes in previous works, such as [8, 11, 16], typically rely on observations of customer delays, where in the proofs it plays a central role that the queue process is geometrically ergodic [8, Lemma 2]. This approach necessitates assuming that the moment-generating and cumulant-generating functions of service and inter-arrival times satisfy certain Cramér-type conditions. Importantly, our price update scheme avoids such assumptions. The process at the center of our analysis is the sequence of inter-arrival times, which is well-behaved (light tails and quick mixing) regardless of the workload process. We assume Poisson arrivals, which is a mild assumption in the context of a large pool of potential customers independently requesting service.
@@
    To establish convergence to the optimal price $p^{*}$, it is necessary at various stages to control the transient behavior of the underlying queueing system. During each cycle of the price-adjustment algorithm, the price is held constant — say at value $p$ — so the system temporarily behaves like an M/G/1 queue with balking customers, where the per-customer joining probability is $H_{\theta}(p,v)$ when the observed congestion level is $v$. Analyzing such queues is notoriously difficult, even in terms of their stationary behavior. We succeed in developing sufficiently sharp bounds, based on a natural coupling idea, that enable proving the convergence to the revenue maximizing price. These enable us to identify explicit bounds on the bias and ‘variability’ of the gradient of the revenue (i.e., the derivative of the revenue with respect to the price, in each step of the algorithm).
@@
    Our approach paves the way for the development of more robust online control algorithms for queueing systems with congestion-dependent joining behavior. Unlike in [8], we do not optimize over service capacity. However, a byproduct of our analysis is the derivation of workload gradients, which makes our method extensible to settings that include holding costs. Notably, these workload gradients differ structurally from those obtained through classical IPA schemes, due to the dependency between inter-arrival times and the workload. Furthermore, our approach naturally extends to multi-server systems, enabling optimization over the number of servers as well. We empirically validate the effectiveness of the resulting price update scheme in the multi-server setting.
@@ 2.1 System dynamics
Note that instead of interpreting $H(\cdot,\cdot)$ as the probability that a customer joins the system, we may view it as follows. Customers are equipped with a
@@
When an incoming customer faces an admission price $p$ and a prospective waiting time $V$, they evaluate their disutility of joining the service, given by $\phi(p,V)$, and decide to enter the system only if $Y\geq\phi(p,V)$; otherwise, they balk. In Example A, we considered the case $\phi(p,V)=\theta_{1}p+\theta_{2}V$ and ${\mathbb{P}}(Y\leq y)=1-e^{-x}$, where the latter corresponds to the disutility threshold $Y$ being exponentially distributed with mean 1. In this paper, for the ease of discussion, we stick with the framework as described originally, i.e., with $H(\cdot,\cdot)$ viewed as a joining probability, and not as a disutility threshold distribution function.
@@
As mentioned, the effective arrival process is not Poisson: the interarrival times are neither exponentially distributed nor independent. In this example we further investigate the distribution of the effective interarrival times.
Figure 2 shows plots for Example A the CDF of an effective interarrival time when $\Lambda=20$ and $H(p,V)=\exp\left(-0.1p-0.2V\right)$. In particular, Figure 2(a) shows plots for a fixed admission price $p=20$, and different initial workloads $\overline{W}\in\{0,5,10,15,20,25,30\}$ whereas Figure 2(b) shows plots for a fixed initial workload $\overline{W}=25$, and different admission prices $p\in\{0,5,10,15,20,25,30\}$.

### v2

@@ Abstract
Keywords. Online revenue maximization $\circ$ Unobserved balking $\circ$ Infinitesimal Perturbation Analysis $\circ$ Regret analysis
SAB is with Korteweg-de Vries Institute for Mathematics, University of Amsterdam, Amsterdam, The Netherlands. HH is with Edwardson School of Industrial Engineering, Purdue University, West Lafayette IN, United States. MM is with the Mathematical Institute, Leiden University, The Netherlands; MM is also affiliated with (a) Korteweg-de Vries Institute for Mathematics, University of Amsterdam, Amsterdam, The Netherlands, b) Eurandom, Eindhoven University of Technology, Eindhoven, The Netherlands, (c) Amsterdam Business School, Faculty of Economics and Business, University of Amsterdam, Amsterdam, The Netherlands. LR is with Department of Statistics, University of Haifa, Israel.
@@ 1 Introduction
In this paper, we examine a service facility modeled as a single-server queue, where customers arrive in response to the price set by the provider. The service provider’s objective is to determine the optimal price, i.e., the price that maximizes the expected revenue per unit of time. Achieving this goal is particularly challenging, as it requires understanding the intricate relationship between pricing, customer demand, system congestion, and the resulting impact on revenue performance.
Offline and online revenue maximization — To put our work into perspective, we first provide a high-level categorization of the existing literature on revenue optimization. Broadly speaking, two dominant approaches can be identified, that one could refer to as the offline and online paradigms.
@@
    Offline revenue optimization is typically studied using a demand curve $d(p)$ that depends on the price $p$. The standard static optimization problem takes the form $\max_{p\geq 0}p\,d(p)$. However, in the context of service systems, an additional layer of complexity arises: congestion can also reduce demand. Specifically, users may decide to balk if the (expected) delay exceeds their individual utility thresholds. In stylized Markovian models, one can explicitly derive the revenue-maximizing price. A canonical example is Naor’s model [22], which features an observable queue and linear waiting costs. In such settings, both the effective demand and the underlying effective arrival rate $\lambda(p)$ depend on the price and the stationary queue-length distribution. The problem of monopolistic pricing in congested systems has been studied extensively under various modeling assumptions; we refer the reader to [13, 7, 12] for comprehensive reviews. In studies where congestion is modeled through a queueing mechanism, the objective function often includes a deduction for queueing-related costs — see, for example, [18, 19], in which a term proportional to the mean stationary queue length is subtracted. These analyses typically assume that the primitives of the underlying queueing model (in particular the service-time distributions) are known.
@@
    In the dynamic pricing literature, the price is iteratively adjusted until, with some level of confidence, it can be said that the revenue-maximizing price has been found. This line of work typically assumes uncertainty in the demand curve, which must be learned, and in addition that the queueing primitives are not, or partly, known. Such a setting naturally calls for online algorithms that learn and optimize simultaneously; we refer to, e.g., [24] for a survey of this strand of the literature. An added layer of complexity arises when customers are delay-sensitive, for instance, when some decide to balk upon observing excessive congestion.
@@
Dynamic revenue maximization problems come in various flavors. First, they may differ in the components included in the revenue function. While the core component reflects income from service, many formulations also include a penalty term for performance degradation — commonly modeled as a function of mean queueing delay. Second, there are differences in the definition of the decision space. In some studies, the objective is solely to determine the optimal price, while others also optimize over additional parameters such as the service rate or the number of servers. Third, there are several variants based on the type of information available to the service provider when making pricing decisions (e.g., the customers’ arrival times, queueing delays, etc.).
@@
Both offline and online approaches have distinct strengths and limitations. We begin by outlining key considerations related to the offline setting. (1) In this setting, it is typically assumed that the revenue can be expressed as a function of price. While evaluating this objective is tractable in simple queueing models, often relying on exponential inter-arrival and service times, it becomes considerably more challenging when incorporating more realistic features, such as general service-time distributions or realistic customer behaviors like balking upon encountering excessive workloads at arrival. (2) The offline approach also requires knowledge of the queueing model’s primitives and the demand curve. While certain customer parameters can often be estimated from historical data (e.g., the service-time distribution), the demand curve is more difficult to infer, as it requires observations across a broad range of pricing scenarios. An important but often overlooked issue is model uncertainty: the effects of estimation errors are typically ignored. (3) When revenue cannot be expressed explicitly, the offline approach frequently resorts to heavy-traffic approximations, which involve scaling up arrival rates and service capacity. This can make the optimal price sensitive to system size, entailing that in contexts where such scaling is inappropriate, the accuracy of these approximations degrades.
@@
We now discuss the advantages and challenges of online pricing approaches. (1) The goal is to develop algorithms with low regret, defined as the cumulative revenue loss relative to the optimal fixed price in hindsight. A key strength of such algorithms is that they operate without requiring an explicit revenue function or prior parameter estimation. Instead, prices are adjusted dynamically based on observed system behavior. (2) The central challenge lies in devising a price update mechanism. A natural strategy is to define cycles within which the price remains fixed, and based on the observed behavior adapt the price. This setting can be framed as a multi-armed bandit problem, wherein each arm corresponds to a different price, and the objective is to identify the price that maximizes expected revenue. However, unlike classical bandits with independent rewards, queueing systems exhibit memory effects: customers carried over from previous cycles influence the current system state and, consequently, the observed revenue. Therefore, any effective algorithm must account for these transient dynamics when evaluating performance at different price levels. (3) Another challenge concerns how prices are updated. A common gradient-based method adjusts the price iteratively using an estimate of the objective’s gradient. The main difficulty is estimating this gradient accurately from data while ensuring fast convergence of the price updating algorithm.
Our setup — As mentioned, the class of revenue maximization problems we consider requires a sound modeling of the interaction between price, demand, and congestion. Existing approaches generally decompose the objective into a demand-based revenue component and a congestion penalty, with the relative influence of each determined by ad hoc weighting parameters. In contrast, our framework integrates congestion effects directly into a single demand-related revenue term, thereby eliminating the need to assign weights to intrinsically incomparable quantities such as monetary revenue and mean queueing delay. Moreover, our model embodies an inherent decentralized decision-making structure, whereby customers independently decide whether to join or balk after observing the congestion level, and the central planner exerts only indirect influence via the admission price.
@@
Our base model is a single-server system with Poisson arrivals and a general service-time distribution. Upon arrival, customers are informed of the current price $p$ set by the administrator and the current workload level $v$, and decide whether to join or balk. Let $H_{\theta}(p,v)$ denote the probability that a customer joins conditional on observing workload $v$ and price $p$, where $\theta$ is a known (possibly multidimensional) parameter characterizing the customer population. This setup is flexible in the sense that it can represent a heterogeneous population of strategic customers with some parametric distribution of customer types that differ in their utility function, or it can simply reflect a behavioral (e.g., empirical) joining rule. The objective of the administrator is to maximize long-term revenue, i.e., they aim to find $p^{*}$ that maximizes
@@
where $V\equiv V(p)$ is the stationary workload at price $p$. Thus, the effective arrival rate in this dynamic pricing problem is
@@
This type of optimization problem naturally lends itself to a stochastic gradient descent (SGD) approach. However, implementing such an approach presents significant challenges: (1) the gradient of $\lambda(p)$ must first be estimated, and (2) the bias and variance of the resulting estimator must be understood to establish meaningful asymptotic guarantees for the algorithm’s performance. Throughout this work, we assume that $\theta$ is known, as it can be efficiently estimated in general settings with balking customers, as discussed in [4, 14].
@@
Online learning for control of queueing systems — Stochastic approximation (SA) algorithms, the class of iterative methods for finding the roots or extrema of functions when only noisy observations are available, constitute a flexible and robust approach for solving stochastic optimization problems [6]); a specific subclass is SGD, which relies on noisy gradient estimates. In the context of queueing systems there is a difficulty that stationary expectations and their gradients are often intractable, which makes direct implementation of standard SA algorithms infeasible. A useful method to overcome this difficulty is infinitesimal perturbation analysis (IPA); see [11] for an extensive introduction to this method and [17] for an example of its use in a queueing context. The regenerative properties of such systems can often be used in order to obtain unbiased estimators of the objective function and to verify convergence to the true solution [10]. In [23] an SA algorithm based on busy periods is used to evaluate the Nash equilibrium for queues with balking customers, and more general strategic customer behavior. Reinforcement learning is another related approach for steady-state optimization of queueing networks [9], particularly well-suited to settings with a complex or high-dimensional policy space. In the context of dynamic pricing of service systems, the analysis is often limited to stylized, tractable models. For example, [1] presents a Bayesian method for learning the optimal admission price in a Markovian system with an unknown proportion of patient customers. The method relies on posterior updating of this proportion based on observed customer decisions.
@@
Our work follows in the tradition of [8], in that it also develops an SGD-based online learning dynamic pricing algorithm aimed at identifying the revenue-maximizing price in a setting that includes congestion effects. In [8], these congestion effects are incorporated into the objective function through a term proportional to the mean stationary queue length, while the arrival rate depends solely on the current price level. It is assumed that the function $\lambda(p)$ is known, while the customers’ service-time characteristics are not. This approach necessitates assuming that the moment-generating and cumulant-generating functions of service and inter-arrival times satisfy certain Cramér-type conditions. The analysis relies on a specific construction for the inter-arrival and service time distributions to enable the use of conventional IPA. Specifically, their setup involves a sequence of independent and identically distributed random variables $(U_{n})_{n\in\mathbb{N}}$, where the $n$-th inter-arrival time is modeled as $U_{n}/\lambda(p)$ given that the price is $p$; a similar construction is applied to the service times. Notably, this definition implies that the effective arrival process is not congestion-dependent. In contrast, our approach avoids imposing such a structure, in that we assume that the inter-arrival distribution depends on both the price and the workload.
@@
    Regarding the problem formulation, we explicitly model the behavior of customers who choose not to join the system when faced with high congestion, resulting in lost revenue. Incorporating such balking customers is particularly challenging, as in our framework we observe only the effective arrivals — that is, data related to the customers who do not balk and enter the system. The dependence of each customer’s joining decision on the congestion level introduces several evident technical hurdles. A useful insight is that the arrival rate at a given price $p$ can be written as the reciprocal of the mean effective inter-arrival time at the same price.
@@
    We develop an online algorithm that provably converges to the optimal price $p^{*}$. The adaptive queue control schemes in previous works, such as [8, 11, 17], typically rely on observations of customer delays, where in the proofs it plays a central role that the queue process is geometrically ergodic [8, Lemma 2], and assumes, as noted before, that the service and inter-arrival times satisfy Cramér-type conditions. Importantly, our price update scheme avoids such assumptions. The process at the center of our analysis is the sequence of inter-arrival times, which is well-behaved (light tails and quick mixing) regardless of the workload process. We assume Poisson arrivals, which is a mild assumption in the context of a large pool of potential customers independently requesting service.
@@
    To establish convergence to the optimal price $p^{*}$, it is necessary at various stages to control the transient behavior of the underlying queueing system. In each cycle of the price-adjustment algorithm, the price is kept fixed at a value $p$. The system then operates as an M/G/1 queue with balking customers, in which the probability that an arriving customer joins is given by $H_{\theta}(p,v)$, conditional on the observed congestion level $v$. Analyzing such queues is notoriously difficult, even in terms of their stationary behavior. We succeed in developing sufficiently sharp bounds, based on a natural coupling idea, that enable proving the convergence to the revenue maximizing price. These enable us to identify explicit bounds on the bias and ‘variability’ of the gradient of the revenue (i.e., the derivative of the revenue with respect to the price, in each step of the algorithm).
@@
    Our approach paves the way for the development of more robust online control algorithms for queueing systems with congestion-dependent joining behavior. Unlike in [8], we do not optimize over service capacity. However, a byproduct of our analysis is the derivation of workload gradients, which makes our algorithm extensible to settings that include holding costs. Notably, these workload gradients differ structurally from those obtained through classical IPA schemes, due to the dependency between inter-arrival times and the workload. Furthermore, our approach naturally extends to multi-server systems, enabling optimization over the number of servers as well.
@@ 2.1 System dynamics
For ease of discussion, throughout our exposition we maintain the probabilistic interpretation of $H$. Nonetheless, there is an alternative (utility-theoretic) interpretation of $H$ as well. Specifically, suppose customers are equipped with
@@
When an incoming customer faces an admission price $p$ and a prospective waiting time $V$, they evaluate their disutility of joining the service, given by $\phi(p,V)$, and decide to enter the system only if $Y\geq\phi(p,V)$; otherwise, they balk. In Example A, we considered the case $\phi(p,V)=\theta_{1}p+\theta_{2}V$ and ${\mathbb{P}}(Y\leq y)=1-e^{-x}$, where the latter corresponds to the disutility threshold $Y$ being exponentially distributed with mean 1.
@@
As mentioned, the effective arrival process is not Poisson: the interarrival times are neither exponentially distributed nor independent. Figure 2 plots the CDF of an effective interarrival time in Example A for $\Lambda=20$, $(\theta_{1},\theta_{2})=(0.1,0.2)$. In particular, Figure 2(a) shows plots for a fixed admission price $p=20$, and different initial workloads $\overline{W}\in\{0,5,10,15,20,25,30\}$ whereas Figure 2(b) shows plots for a fixed initial workload $\overline{W}=25$, and different admission prices $p\in\{0,5,10,15,20,25,30\}$.